When one half of the second of two consecutive integers is subtracted from the first of the two consecutive integers, the result is 2.
Answer by bucky(2189)

Since the two integers are consecutive, if we call the first one X, then the second one
will be 1 number bigger, so it will be X + 1.
.
One-half of the second integer is (1/2)*(X + 1). If you subtract this from the first integer
X the answer is to be 2. In equation form this is:
.
X - (1/2)(X + 1) = 2
.
You can get rid of the fraction (1/2) by multiplying both sides of this equation (all terms) by
2 and the equation becomes:
.
2X - (X + 1) = 4
.
Since the parentheses are preceded by a minus sign, you can remove the parentheses if
you change the signs of the terms inside the parentheses. Applying this rule results in:
.
2X - X - 1 = 4
.
Get rid of the -1 on the left side by adding +1 to both sides and you have:
.
2X - X = 5
.
Combine the two terms on the left side and you have:
.
X = 5
.
This tells you that the two consecutive integers are 5 and 6
.
Check. If you take half of the second integer you get 3. And if you subtract that from
the first integer you get 5 - 3 = 2 ... which is just what the problem required. Therefore,
the answer that the integers are 5 and 6 checks out.
.
Hope this helps you to understand how to work the problem.
.